OpenStudy (anonymous):

cosxsin2x/1+cos2x simplify

OpenStudy (zehanz):

You probably mean:\[\frac{ \cos x \sin 2x }{ 1+\cos 2x }\] To simplify, you need to apply the double-angle formulas: sin2x=2sinxcosx cos2x=1-2sin²x=2cos²x-1=cos²x-sin²x Now can you honestly tell me you've learned these very important formulas by heart?

OpenStudy (anonymous):

nope not yet i just started with these

OpenStudy (zehanz):

OK, so learn them ;) Now replace sin2x with 2sinxcosx and cos2x with one of the co2x identities. See if you can simplify after this.

OpenStudy (anonymous):

2cos^2x= 1 + cos 2x---> numerator = 2cos^2x. denominator cosx *sin2x = 2sinxcosx . *cosx= 2cos^2x sinx . combine both to get sinx. is it help?

OpenStudy (zehanz):

If I write it in the equation editor, it is much clearer for me:\[\frac{ \cos x 2\sin x \cos x }{ 1+2\cos^2x-1 }=\frac{ 2\sin x \cos^2x }{ 2\cos^2x }=\]I think you'll see where this is going to...

OpenStudy (anonymous):

i have no idea of what the next step is

OpenStudy (anonymous):

cancel out the same term from both sides

OpenStudy (anonymous):

so then ill stay with sinx?

OpenStudy (zehanz):

Yup!

OpenStudy (anonymous):

thats my biggest problem. idk how to do the correct steps.

OpenStudy (zehanz):

Practising makes that (gradually) better, believe me!

OpenStudy (zehanz):

Next time, chances are you can get one step further yourself...

OpenStudy (anonymous):

Thank you for ur help, i'll post another one see if u can guide me through it

OpenStudy (zehanz):

YW!